# Neutrinos, B-L Symmetry and the Dark Dimension

### v1

@@ 1 Introduction
The Dark Dimension scenario Montero:2022prj has been motivated from Swampland principles, and in particular from the Distance Conjecture Ooguri:2006in; Lust:2019zwm applied to Dark Energy. In a nutshell, it proposes the existence of one mesoscopic extra dimension in our universe with diameter in the $0.1-10\ \mu m$ range, which is significantly larger than other extra dimensions. This large dimension (dubbed the Dark Dimension) is motivated from the smallness of dark energy $\Lambda$ and leads to the existence of a light infinite tower of states whose mass is correlated with the value of the dark energy. Theoretical and experimental bounds single out this tower to be a Kaluza-Klein tower of mass $m_{\rm KK}\sim\Lambda^{1/4}$, signaling decompactification of only one extra dimension Montero:2022prj [^1] See Anchordoqui:2025nmb; Hardy:2025ajb for a recent discussion on the borderline experimental feasability of having two large dark dimensions if the diameter is smaller $\lesssim 0.3\ \mu m$.. In this scenario, the SM fields should live in a brane localised in the Dark Dimension. Although it remains as an open question why the Dark Energy is so small in the first place, it provides a quantum gravity link between the Dark Energy and diverse aspects of Particle Physics.
@@
This has led to a very rich phenomenology addressing open puzzles in Particle Physics and Cosmology. For instance, the Dark Dimension scenario has led to new models for dark matter and cosmology Anchordoqui:2022txe; Gonzalo:2022jac; Anchordoqui:2022tgp; Anchordoqui:2022svl; Anchordoqui:2023oqm; Noble:2023mfw; Anchordoqui:2023wkm; Anchordoqui:2023tln; Law-Smith:2023czn; Cui:2023wzo; Obied:2023clp; Anchordoqui:2024akj; Anchordoqui:2024dxu; Gendler:2024gdo; Anchordoqui:2024gfa; Heckman:2024trz; Basile:2024lcz; Anchordoqui:2024xvl; Anchordoqui:2025xug; Antoniadis:2025rck; Eller:2025lsh. In particular Bedroya:2025fwh, based on the earlier model Agrawal:2019dlm, leads to the best known fit of a physical model to the recent cosmological observations of evolving dark sector DESI:2025zgx; DESI:2025fii. The Dark Dimension scenario also has a distinct experimental prediction for deviations from Newton’s inverse square law at micron length scales (see Adelberger:2003zx; ParticleDataGroup:2022pth; Lee:2020zjt for reviews) which is currently being pursued by a number of groups.
@@
It is thus natural to ask how various Beyond Standard Model (BSM) questions fit in this scenario. For example, see Heckman:2024trz for a discussion of how GUTs may work with this scenario. Another glaring question which demands a BSM explanation is the existence of mass for neutrinos. A proposal of how neutrinos may fit was already presented in Montero:2022prj, along the lines of similar discussions in Large Extra Dimension scenarios, see Arkani-Hamed:1998wuz; Dienes:1998sb, by allowing massless bulk five-dimensional right-handed neutrinos (see Antoniadis:2025rck for a recent study suggesting these may be within experimental reach). As we will review below, this proposal has the benefit that it utilizes (but does not explain from fundamental principles) the observed coincidence between the mass scale for neutrinos and that of Dark Energy. On the other hand, postulating that there are at least three massless fermions in the bulk just to play the role of right-handed neutrino and provide mass to the neutrinos seems somewhat ad hoc.
@@ 2 Neutrino masses in the Dark Dimension
We begin by reviewing the simplest ways to generate neutrino masses in the context of the Dark Dimension scenario. Of course, the simplest possibility is that the SM fields embed in a standard GUT scenario with a gauge group like $SO(10)$, where the right-handed neutrinos are unified with the rest of the SM fields. Such a model is completely compatible with the Dark Dimension despite the fact that GUT scale is far above the 5d Planck scale Heckman:2024trz, as the heavy GUT states are realized as stretched strings. But here, we would like to explore also other options which may be even more natural in the context of the dark dimensions scenario.
In the Dark Dimension (DD) scenario, the SM fields live in a localized brane. After EW symmetry breaking, this contains massless left-handed neutrinos, for which masses must be generated. One simple possibility, and quite natural from an extra-dimensional point of view, arises by assuming the existence of massless, five-dimensional bulk fermions. In four-dimensional language, these generate a tower of sterile, right-handed neutrinos with mass scale that can be naturally of the order of tower scale $m_{\nu_{s}}\sim\Lambda^{1/4}$, which roughly coincides with the neutrino mass scale Montero:2022prj. The idea that bulk fermions can play such a role was also considered in the Large extra dimension models Dienes:1998sb; Arkani-Hamed:1998wuz, though in that context one needs to assume unnaturally one dimension is much larger than the rest, whereas this is automatic for the Dark Dimension scenario. In the DD scenrio, one expects the light left-handed neutrinos on the GUT brane to couple to the KK tower of right-handed ones in the bulk (see e.g. Carena:2017qhd). Schematically this leads, for each flavor, to an effective $2\times 2$ mass matrix of the form
@@
where $R\sim 1/m_{KK}$ is the size of the extra dimension, $M_{5}\sim\Lambda^{1/12}\sim 10^{10}$ GeV is the 5d Planck scale, $\langle H\rangle$ is the Higgs vev and $y$ the Yukawa coupling of the SM brane with the bulk states. The $1/{\sqrt{R}}$ comes from the wave function normalization of the bulk 5d fermions, assuming a homogeneous space, and we have replaced the infinite tower of right-handed modes by a single one for simplicity, as it does not change the qualitative features. See Montero:2022prj; Dienes:1998sb for details in the general case.
@@
modulo $\mathcal{O}(1)$ factors that are controlled by the effective number of states in the tower, and can be tuned by turning on a bulk mass for the neutrino, if necessary. This is an expression similar to the usual see-saw mechanism with a Yukawa coupling of order $y\sim 10^{-3}$. The rest of the mass eigenstates will have an increasing mass starting at $m_{KK}$. See Antoniadis:2025rck; Eller:2025lsh for recent analysis on the experimental signatures of these bulk right-handed neutrinos in the context of the Dark Dimension scenario.
Having $m_{s}\gtrsim\frac{y\langle H\rangle}{\sqrt{lM_{5}}}$ is essential to obtain the correct value of the neutrino masses, and this hierarchy demands an explanation. This was addressed in Gonzalo:2022jac, by imposing that the moduli of the SM brane has had enough time during the cosmological evolution of the universe to decay away, which leads to their mass being $\gtrsim\Lambda^{1/6}$. If we take this as the ballpark for $H$, it leads to $m_{\nu}/y^{2}\gtrsim\Lambda^{2/6}/\Lambda^{1/12}\sim\Lambda^{1/4}$, leading to a potential explanation of the hierarchy between $m_{s}\sim\Lambda^{1/4}$ and the off-diagonal terms. Even though the parameter range works very well, modulo the somewhat small value for the Yukawa coupling $y$, it should be emphasized that the bound $m_{SM}\gtrsim\Lambda^{1/6}$ is not a built-in feature of the Dark Dimension scenario – rather, it is imposed on the actual model in order to realize a sensible cosmology compatible with observations. Therefore the observed coincidence $m_{\nu}\sim m_{KK}$ is anthropic in this model – it is only there because, if it was not, many scalars would not have stabilized by now, perhaps preventing the emergence of observers such as ourselves. By contrast, the coincidence $m_{KK}\sim\Lambda^{1/4}$ is a fundamental feature of the Dark Dimension scenario, irrespectively of other coincidences.
@@
This implies that each bulk fermion only yields effectively a single lower-dimensional mode that couples to $\nu_{L}$, so that we cannot generate masses for the three SM left-handed neutrinos if starting with only one bulk fermion. We need to have at least three bulk fermionic fields to do so. In contrast, if we give the SM brane some size, then a massive mode of the bulk fermion which has wavelength as the size of the SM brane may give an extra mode. But the mass of such a mode can be estimated to be roughly $1/L$, where $L$ is the SM brane size. At this scale, we also expect to find the KK copies of the rest of the SM fields. The fact that these have not been observed by LHC means that this mass scale has to be at least 10 TeV (and is expected to be of that order in the GUT realization of dark dimensions Heckman:2024trz). Such a large value for $m_{KK}$ would lead, due to (2.1), to a neutrino mass
@@ 3 The bulk B-L scenario
In principle, we could take the Dark Dimension to be an interval $S^{1}/\mathbb{Z}_{2}$ (as in a Horava-Witten setup Horava:1995qa; Horava:1996ma or an orbifold in string theory). This has been discussed in the context of Dark Dimension in Schwarz:2024tet (see also Reig:2025dpz), though what we have to say in this paper is not restricted to this case. If we take the fifth dimension to be an interval, the question arises as to what are the boundary conditions of the 5d gauge field $A_{B}$ at the endpoints. While there are infinitely many possibilities Arbalestrier:2025jsg, we will focus here on two simple ones:
@@ 3.1 The pseudovector case
As we will now see, the pseudovector case (where the longitudinal component has Dirichlet boundary conditions) leads to no bulk neutrinos – instead, B-L anomaly cancellation happens via anomaly inflow from the bulk, very much like in an $E_{8}$ Horava-Witten wall in string theory Horava:1995qa; Horava:1996ma. The similitude arises from the fact that the three-form $C_{3}$ in M-theory transforms as a pseudo-three-form under parity. In the four-dimensional language, the anomaly inflow leads to a Green-Schwarz mechanism Green:2012pqa.
@@
Notice that what we have here is slightly different from standard anomaly inflow of the kind we have e.g. on heterotic strings. There, an anomaly of a global worldsheet symmetry coupled to a bulk gauge field is cancelled by a bulk Chern-Simons coupling. Here, we use the bulk to cancel the anomaly of a worldvolume gauge field, that mixes with the bulk. One could say that $A_{X}$ is just the restriction of $A_{B}$ to the brane, but they don’t have to agree, just their fieldstrengths have as in (3.8). In the HW case they are clearly different fields, as an $E_{8}$ gauge field cannot be regarded as a restriction of $G_{4}$ to a boundary (only at the topological level). In modern language, one identifies boundary conditions for the $U(1)$ gauge field $A_{B}$ with quantum field theories living in the boundary and equipped with a $U(1)_{B}$ global symmetry Arbalestrier:2025jsg; in this case, the Chern-Simons terms of the bulk theory specify the mixed gravitational-$U(1)_{B}$ anomalies of the boundary theory.
@@
On an interval compactification, $A_{B}^{(4)}$ is projected out, since the existence of Chern-Simons terms (3.9) force again parity transformations to act with an additional minus sign on $A_{B}$ (again, just like in M-theory, where $C_{3}$ gets an additional - sign for similar reasons Witten:1996md). The $A_{X}$ field survives directly in 4d, and the BF coupling (3.8) can be dualized to give a Stuckelberg term
@@
It is important to note that, for the above mechanism to work, the anomaly polynomial of $U(1)_{X}$ must be severely constrained. It must be of the form (3.9), i.e. it can have pure $U(1)_{X}^{3}$ and mixed $U(1)_{X}$-gravitational anomalies, but it cannot have any mixed anomalies with SM gauge fields. This is similar to how the Horava-Witten factorization only works with very special groups Horava:1996ma where all anomalies can be written purely in terms of $c_{2}$. In other words, this mechanism explains why B-L is a good approximate symmetry of the SM, and why it has no mixed anomalies with it, but it does have pure and mixed gravitational ones – these are precisely the ones that can be cancelled by the Green-Schwarz mechanism. One would have expected that a random global symmetry of the SM would have had mixed anomalies with SM gauge fields. The fact that B-L does not suggests it has an ulterior meaning, and the above mechanism is a scenario where this can be explained naturally without having any new particles. In this sense, it is similar to the mechanisms in Wang:2020mra; Wang:2025oow. Notice that, in this case, bulk massless particles charged under $U(1)_{B}$ cannot alter the anomaly, since their zero modes are not charged under the gauge fields.
@@ 3.2 The vector case
where the choice $\pm$ depends on the intrinsic parity of the 5d fermion. Therefore, one of the two chiralities is projected out, and we end up with a single four-dimensional Weyl fermion of charge qq [^3] An $U(1)_{B}$-invariant 5d mass term for the fermion is forbidden by parity as we will see below, which is why the massless four-dimensional zero mode is robust. This is anomalous, so a suitable collection of 5d fermions can cancel the anomaly of the SM gauge fields. The simplest choice is to have three massless bulk fermions, each of charge one, which upon dimensional reduction yield three four-dimensional right-handed neutrinos, ensuring anomaly cancellation of BLB-L. Therefore, in this scenario, the presence of three massless bulk neutrinos is naturally explained. Anomaly cancellation between bulk fields and localized sectors (twisted sectors in orbifolds) is a familiar phenomenon in string theory, see e.g. Dixon:1985jw for an early realization.
@@
In Section 2 and in Gonzalo:2022jac the same coincidence was explained from the condition that the moduli controling SM brane has to have at least a mass of order $\Lambda^{1/6}$ to have had time to decay in Hubble time, and this in turn naturally sets the scale for Higgs potential, together with a tuning of the relevant Yukawas by three orders of magnitude. By contrast, (3.26) explains the coincidence between neutrino mass scale and cosmological constant much more naturally, provided that the vevs of $\Phi$ and the Higgs field are related. Hence, it is precisely the lack of fine-tunings in this scenario that explains this intriguing coincidence between neutrino masses and dark energy.
@@
This is in contrast with the result $m_{\nu_{s}}\sim\Lambda^{1/4}$ of Montero:2022prj summarized in Section 2, which led to fine-tuning the Yukawa coupling $y$ to get the correct mass for active neutrinos. It would be important to look for experimental signatures of sterile neutrinos in this mass range.
@@
Constraints on the $(m_{B-L},g_{B-L})$ parameter space can be found e.g. in Heeck:2014zfa; Ilten:2018crw; Escudero:2018fwn; Herbermann:2025uqz; Asai:2023mzl; Wang:2024gvt and depicted in Figure 1. For mases around 100 GeV, the dominant constraints come from LHC searches. The most natural value from the Dark Dimension perspective, $g_{B-L}\sim 10^{-10}$ and $m_{B-L}\approx 100$ GeV is safely within this allowed region. One might worry that, since $m_{B-L}\sim\,100$ GeV, future LHC searches (which see up to $\approx 100$ times this energy) may also see the higher KK modes of the BLB-L vector boson, thereby dramatically increasing the production cross-section. If one has $N$ almost degenerate copies of the vector participating, a tree-level cross-section scales like
@@
At $s\gg m_{B-L}^{2}$, we recognize the linear running of $g_{B-L}^{2}$ with energy characteristic of 5d gauge theory. Since $m_{KK}$ is so small, even at energies $\sqrt{s}$ slightly above $m_{B-L}$ we will get a significant enhancement of the effective gauge coupling. For $\alpha\sim 1$, setting $\sqrt{s}\sim 10$ TeV leads to $g_{B-L,\text{effective}}\sim 10^{-3}$. While this is indeed much larger than our naive estimate, it is still within the allowed bounds, but perhaps within experimental reach of future colliders. At any rate, the analysis we just carried out should be regarded as a first step only, since the precise effects of the KK towers of both B-L and right-handed bulk neutrino fields will depend on the details of the process under consideration (for instance, for LEP bounds, one should use the LEP $\sqrt{s}\sim 100$ GeV, which is below B-L gauge boson mass, in which case again $g_{B-L}\sim 10^{-10}$ is very far from the experimental bounds). Astrophysical bounds on B-L are not relevant since our B-L gauge boson is too heavy to be produced in supernovae Esseili:2023ldf, and neither are BBN bounds for the same reason. There can be strong bounds if the right-handed neutrinos are produced during the cosmological evolution of the universe Herbermann:2025uqz, but if the initial temperature of Gonzalo:2022jac for the SM fields is used, then the bounds are not relevant either. We conclude that this scenario is predictive, and quite appealing. If LHC bounds on $g_{B-L}$ improve significantly, it might be possible to put constraints on various parameters of the model.
Figure 1: Theoretical prediction versus experimental constraints for the gauge coupling and mass of the BLB-L gauge boson. Our benchmark model predicted value is shown as a blue star, with a shaded band indicating the uncertainty associated with the precise value of $\alpha$ and the energy at which the gauge coupling is evaluated. We also show experimental limits from BABAR BaBar:2017tiz, LHCb LHCb:2017trq, Drell-Yan process in ATLAS Escudero:2018fwn; ATLAS:2017rue ATLAS ATLAS:2017fih, LEP (taken from Asai:2023mzl; Wang:2024gvt), and BBN limits taken from Heeck:2014zfa; Herbermann:2025uqz.

### v2

@@ 1 Introduction
The Dark Dimension scenario Montero et al. (2023) has been motivated from Swampland principles, and in particular from the Distance Conjecture Ooguri and Vafa (2007); Lüst et al. (2019) applied to Dark Energy. In a nutshell, it proposes the existence of one mesoscopic extra dimension in our universe with diameter in the $0.1-10\ \mu m$ range, which is significantly larger than other extra dimensions. This large dimension (dubbed the Dark Dimension) is motivated from the smallness of dark energy $\Lambda$ and leads to the existence of a light infinite tower of states whose mass is correlated with the value of the dark energy. Theoretical and experimental bounds single out this tower to be a Kaluza-Klein tower of mass $m_{\rm KK}\sim\Lambda^{1/4}$, signaling decompactification of only one extra dimension Montero et al. (2023) [^1] See Anchordoqui et al. (2025b); Hardy et al. (2025) for a recent discussion on the borderline experimental feasability of having two large dark dimensions if the diameter is smaller $\lesssim 0.3\ \mu m$.. In this scenario, the SM fields should live in a brane localised in the Dark Dimension. Although it remains as an open question why the Dark Energy is so small in the first place, it provides a quantum gravity link between the Dark Energy and diverse aspects of Particle Physics.
@@
This has led to a very rich phenomenology addressing open puzzles in Particle Physics and Cosmology. For instance, the Dark Dimension scenario has led to new models for dark matter and cosmology Anchordoqui et al. (2022); Gonzalo et al. (2023); Anchordoqui et al. (2023c, b, a); Noble et al. (2023); Anchordoqui et al. (2024a, d); Law-Smith et al. (2024); Cui and Ning (2023); Obied et al. (2024); Anchordoqui et al. (2024c, e); Gendler and Vafa (2024); Anchordoqui et al. (2024b); Heckman et al. (2025); Basile and Lust (2025); Anchordoqui et al. (2025a, c); Antoniadis et al. (2025); Eller et al. (2025). In particular Bedroya et al. (2025), based on the earlier model Agrawal et al. (2021), leads to the best known fit of a physical model to the recent cosmological observations of evolving dark sector Abdul Karim and others (2025); Lodha and others (2025). The Dark Dimension scenario also has a distinct experimental prediction for deviations from Newton’s inverse square law at micron length scales (see Adelberger et al. (2003); Workman and others (2022); Lee et al. (2020) for reviews) which is currently being pursued by a number of groups.
@@
It is thus natural to ask how various Beyond Standard Model (BSM) questions fit in this scenario. For example, see Heckman et al. (2025) for a discussion of how GUTs may work with this scenario. Another glaring question which demands a BSM explanation is the existence of mass for neutrinos. A proposal of how neutrinos may fit was already presented in Montero et al. (2023), along the lines of similar discussions in Large Extra Dimension scenarios, see Arkani-Hamed et al. (2001); Dienes et al. (1999), by allowing massless bulk five-dimensional right-handed neutrinos (see Antoniadis et al. (2025) for a recent study suggesting these may be within experimental reach). As we will review below, this proposal has the benefit that it utilizes (but does not explain from fundamental principles) the observed coincidence between the mass scale for neutrinos and that of Dark Energy. On the other hand, postulating that there are at least three massless fermions in the bulk just to play the role of right-handed neutrino and provide mass to the neutrinos seems somewhat ad hoc.
@@ 2 Neutrino masses in the Dark Dimension
We begin by reviewing the simplest ways to generate neutrino masses in the context of the Dark Dimension scenario. Of course, the simplest possibility is that the SM fields embed in a standard GUT scenario with a gauge group like $SO(10)$, where the right-handed neutrinos are unified with the rest of the SM fields. Such a model is completely compatible with the Dark Dimension despite the fact that GUT scale is far above the 5d Planck scale Heckman et al. (2025), as the heavy GUT states are realized as stretched strings. But here, we would like to explore also other options which may be even more natural in the context of the dark dimensions scenario.
In the Dark Dimension (DD) scenario, the SM fields live in a localized brane. After EW symmetry breaking, this contains massless left-handed neutrinos, for which masses must be generated. One simple possibility, and quite natural from an extra-dimensional point of view, arises by assuming the existence of massless, five-dimensional bulk fermions. In four-dimensional language, these generate a tower of sterile, right-handed neutrinos with mass scale that can be naturally of the order of tower scale $m_{\nu_{s}}\sim\Lambda^{1/4}$, which roughly coincides with the neutrino mass scale Montero et al. (2023). The idea that bulk fermions can play such a role was also considered in the Large extra dimension models Dienes et al. (1999); Arkani-Hamed et al. (2001), though in that context one needs to assume unnaturally one dimension is much larger than the rest, whereas this is automatic for the Dark Dimension scenario. In the DD scenrio, one expects the light left-handed neutrinos on the GUT brane to couple to the KK tower of right-handed ones in the bulk (see e.g. Carena et al. (2017)). Schematically this leads, for each flavor, to an effective $2\times 2$ mass matrix of the form
@@
where $R\sim 1/m_{KK}$ is the size of the extra dimension, $M_{5}\sim\Lambda^{1/12}\sim 10^{10}$ GeV is the 5d Planck scale, $\langle H\rangle$ is the Higgs vev and $y$ the Yukawa coupling of the SM brane with the bulk states. The $1/{\sqrt{R}}$ comes from the wave function normalization of the bulk 5d fermions, assuming a homogeneous space, and we have replaced the infinite tower of right-handed modes by a single one for simplicity, as it does not change the qualitative features. See Montero et al. (2023); Dienes et al. (1999) for details in the general case.
@@
modulo $\mathcal{O}(1)$ factors that are controlled by the effective number of states in the tower, and can be tuned by turning on a bulk mass for the neutrino, if necessary. This is an expression similar to the usual see-saw mechanism with a Yukawa coupling of order $y\sim 10^{-3}$. The rest of the mass eigenstates will have an increasing mass starting at $m_{KK}$. See Antoniadis et al. (2025); Eller et al. (2025) for recent analysis on the experimental signatures of these bulk right-handed neutrinos in the context of the Dark Dimension scenario.
Having $m_{s}\gtrsim\frac{y\langle H\rangle}{\sqrt{lM_{5}}}$ is essential to obtain the correct value of the neutrino masses, and this hierarchy demands an explanation. This was addressed in Gonzalo et al. (2023), by imposing that the moduli of the SM brane has had enough time during the cosmological evolution of the universe to decay away, which leads to their mass being $\gtrsim\Lambda^{1/6}$. If we take this as the ballpark for $H$, it leads to $m_{\nu}/y^{2}\gtrsim\Lambda^{2/6}/\Lambda^{1/12}\sim\Lambda^{1/4}$, leading to a potential explanation of the hierarchy between $m_{s}\sim\Lambda^{1/4}$ and the off-diagonal terms. Even though the parameter range works very well, modulo the somewhat small value for the Yukawa coupling $y$, it should be emphasized that the bound $m_{SM}\gtrsim\Lambda^{1/6}$ is not a built-in feature of the Dark Dimension scenario – rather, it is imposed on the actual model in order to realize a sensible cosmology compatible with observations. Therefore the observed coincidence $m_{\nu}\sim m_{KK}$ is anthropic in this model – it is only there because, if it was not, many scalars would not have stabilized by now, perhaps preventing the emergence of observers such as ourselves. By contrast, the coincidence $m_{KK}\sim\Lambda^{1/4}$ is a fundamental feature of the Dark Dimension scenario, irrespectively of other coincidences.
@@
This implies that each bulk fermion only yields effectively a single lower-dimensional mode that couples to $\nu_{L}$, so that we cannot generate masses for the three SM left-handed neutrinos if starting with only one bulk fermion. We need to have at least three bulk fermionic fields to do so. In contrast, if we give the SM brane some size, then a massive mode of the bulk fermion which has wavelength as the size of the SM brane may give an extra mode. But the mass of such a mode can be estimated to be roughly $1/L$, where $L$ is the SM brane size. At this scale, we also expect to find the KK copies of the rest of the SM fields. The fact that these have not been observed by LHC means that this mass scale has to be at least 10 TeV (and is expected to be of that order in the GUT realization of dark dimensions Heckman et al. (2025)). Such a large value for $m_{KK}$ would lead, due to (2.1), to a neutrino mass
@@ 3 The bulk B-L scenario
In principle, we could take the Dark Dimension to be an interval $S^{1}/\mathbb{Z}_{2}$ (as in a Horava-Witten setup Horava and Witten (1996b, a) or an orbifold in string theory). This has been discussed in the context of Dark Dimension in Schwarz (2024) (see also Reig and Ruiz (2025)), though what we have to say in this paper is not restricted to this case. If we take the fifth dimension to be an interval, the question arises as to what are the boundary conditions of the 5d gauge field $A_{B}$ at the endpoints. While there are infinitely many possibilities Arbalestrier et al. (2025), we will focus here on two simple ones:
@@ 3.1 The pseudovector case
As we will now see, the pseudovector case (where the longitudinal component has Dirichlet boundary conditions) leads to no bulk neutrinos – instead, B-L anomaly cancellation happens via anomaly inflow from the bulk, very much like in an $E_{8}$ Horava-Witten wall in string theory Horava and Witten (1996b, a). The similitude arises from the fact that the three-form $C_{3}$ in M-theory transforms as a pseudo-three-form under parity. In the four-dimensional language, the anomaly inflow leads to a Green-Schwarz mechanism Green et al. (2012).
@@
Notice that what we have here is slightly different from standard anomaly inflow of the kind we have e.g. on heterotic strings. There, an anomaly of a global worldsheet symmetry coupled to a bulk gauge field is cancelled by a bulk Chern-Simons coupling. Here, we use the bulk to cancel the anomaly of a worldvolume gauge field, that mixes with the bulk. One could say that $A_{X}$ is just the restriction of $A_{B}$ to the brane, but they don’t have to agree, just their fieldstrengths have as in (3.8). In the HW case they are clearly different fields, as an $E_{8}$ gauge field cannot be regarded as a restriction of $G_{4}$ to a boundary (only at the topological level). In modern language, one identifies boundary conditions for the $U(1)$ gauge field $A_{B}$ with quantum field theories living in the boundary and equipped with a $U(1)_{B}$ global symmetry Arbalestrier et al. (2025); in this case, the Chern-Simons terms of the bulk theory specify the mixed gravitational-$U(1)_{B}$ anomalies of the boundary theory.
@@
On an interval compactification, $A_{B}^{(4)}$ is projected out, since the existence of Chern-Simons terms (3.9) force again parity transformations to act with an additional minus sign on $A_{B}$ (again, just like in M-theory, where $C_{3}$ gets an additional - sign for similar reasons Witten (1997)). The $A_{X}$ field survives directly in 4d, and the BF coupling (3.8) can be dualized to give a Stuckelberg term
@@
It is important to note that, for the above mechanism to work, the anomaly polynomial of $U(1)_{X}$ must be severely constrained. It must be of the form (3.9), i.e. it can have pure $U(1)_{X}^{3}$ and mixed $U(1)_{X}$-gravitational anomalies, but it cannot have any mixed anomalies with SM gauge fields. This is similar to how the Horava-Witten factorization only works with very special groups Horava and Witten (1996a) where all anomalies can be written purely in terms of $c_{2}$. In other words, this mechanism explains why B-L is a good approximate symmetry of the SM, and why it has no mixed anomalies with it, but it does have pure and mixed gravitational ones – these are precisely the ones that can be cancelled by the Green-Schwarz mechanism. One would have expected that a random global symmetry of the SM would have had mixed anomalies with SM gauge fields. The fact that B-L does not suggests it has an ulterior meaning, and the above mechanism is a scenario where this can be explained naturally without having any new particles. In this sense, it is similar to the mechanisms in Wang (2021, 2025). Notice that, in this case, bulk massless particles charged under $U(1)_{B}$ cannot alter the anomaly, since their zero modes are not charged under the gauge fields.
@@ 3.2 The vector case
where the choice $\pm$ depends on the intrinsic parity of the 5d fermion. Therefore, one of the two chiralities is projected out, and we end up with a single four-dimensional Weyl fermion of charge qq [^3] An $U(1)_{B}$-invariant 5d mass term for the fermion is forbidden by parity as we will see below, which is why the massless four-dimensional zero mode is robust. This is anomalous, so a suitable collection of 5d fermions can cancel the anomaly of the SM gauge fields. The simplest choice is to have three massless bulk fermions, each of charge one, which upon dimensional reduction yield three four-dimensional right-handed neutrinos, ensuring anomaly cancellation of BLB-L. Therefore, in this scenario, the presence of three massless bulk neutrinos is naturally explained. Anomaly cancellation between bulk fields and localized sectors (twisted sectors in orbifolds) is a familiar phenomenon in string theory, see e.g. Dixon et al. (1985) for an early realization.
@@
In Section 2 and in Gonzalo et al. (2023) the same coincidence was explained from the condition that the moduli controling SM brane has to have at least a mass of order $\Lambda^{1/6}$ to have had time to decay in Hubble time, and this in turn naturally sets the scale for Higgs potential, together with a tuning of the relevant Yukawas by three orders of magnitude. By contrast, (3.26) explains the coincidence between neutrino mass scale and cosmological constant much more naturally, provided that the vevs of $\Phi$ and the Higgs field are related. Hence, it is precisely the lack of fine-tunings in this scenario that explains this intriguing coincidence between neutrino masses and dark energy.
@@
This is in contrast with the result $m_{\nu_{s}}\sim\Lambda^{1/4}$ of Montero et al. (2023) summarized in Section 2, which led to fine-tuning the Yukawa coupling $y$ to get the correct mass for active neutrinos. It would be important to look for experimental signatures of sterile neutrinos in this mass range.
@@
Constraints on the $(m_{B-L},g_{B-L})$ parameter space can be found e.g. in Heeck (2014); Ilten et al. (2018); Escudero et al. (2018); Herbermann and Lindner (2025); Asai et al. (2024); Wang et al. (2025) and depicted in Figure 1. For mases around 100 GeV, the dominant constraints come from LHC searches. The most natural value from the Dark Dimension perspective, $g_{B-L}\sim 10^{-10}$ and $m_{B-L}\approx 100$ GeV is safely within this allowed region. One might worry that, since $m_{B-L}\sim\,100$ GeV, future LHC searches (which see up to $\approx 100$ times this energy) may also see the higher KK modes of the BLB-L vector boson, thereby dramatically increasing the production cross-section. If one has $N$ almost degenerate copies of the vector participating, a tree-level cross-section scales like
@@
At $s\gg m_{B-L}^{2}$, we recognize the linear running of $g_{B-L}^{2}$ with energy characteristic of 5d gauge theory. Since $m_{KK}$ is so small, even at energies $\sqrt{s}$ slightly above $m_{B-L}$ we will get a significant enhancement of the effective gauge coupling. For $\alpha\sim 1$, setting $\sqrt{s}\sim 10$ TeV leads to $g_{B-L,\text{effective}}\sim 10^{-3}$. While this is indeed much larger than our naive estimate, it is still within the allowed bounds, but perhaps within experimental reach of future colliders. At any rate, the analysis we just carried out should be regarded as a first step only, since the precise effects of the KK towers of both B-L and right-handed bulk neutrino fields will depend on the details of the process under consideration (for instance, for LEP bounds, one should use the LEP $\sqrt{s}\sim 100$ GeV, which is below B-L gauge boson mass, in which case again $g_{B-L}\sim 10^{-10}$ is very far from the experimental bounds). Astrophysical bounds on B-L are not relevant since our B-L gauge boson is too heavy to be produced in supernovae Esseili and Kribs (2024), and neither are BBN bounds for the same reason. There can be strong bounds if the right-handed neutrinos are produced during the cosmological evolution of the universe Herbermann and Lindner (2025), but if the initial temperature of Gonzalo et al. (2023) for the SM fields is used, then the bounds are not relevant either. We conclude that this scenario is predictive, and quite appealing. If LHC bounds on $g_{B-L}$ improve significantly, it might be possible to put constraints on various parameters of the model.
Figure 1: Theoretical prediction versus experimental constraints for the gauge coupling and mass of the BLB-L gauge boson. Our benchmark model predicted value is shown as a blue star, with a shaded band indicating the uncertainty associated with the precise value of $\alpha$ and the energy at which the gauge coupling is evaluated. We also show experimental limits from BABAR Lees and others (2017), LHCb Aaij and others (2018), Drell-Yan process in ATLAS Escudero et al. (2018); Aaboud and others (2017a) ATLAS Aaboud and others (2017b), LEP (taken from Asai et al. (2024); Wang et al. (2025)), and BBN limits taken from Heeck (2014); Herbermann and Lindner (2025).